Shift report 20011101 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Nov 01 2001 - 15:55:14 EST

  • Next message: Apache: "Shift report 20011102 16:00-24:00"

    ------------------------------------------------------------
    Date: Thu Nov  1 15:55:14 2001, Shift: 08:00-16:00
    Supervisor: BB, with: none
    Log book#10 pages 82 - 84
    
    Shift Summary:
    ------------------------------------------------------------
    1.Shift started out with 2 good runs  ZDC ~470 Hz
    
    2.	Run 5616 with 101k trigger6
    	Run 5617 with 17k trigger6
    Although no changes to run page or parameters were
     made between runs, run 5616 was not written to HPSS
     while run 5617 was.
    
    Beam died just after run 5618 was started.
    
    No beam for the rest of the shift,"Studies for Phenix 
    	beta ramp".
    
    All magnet polarities were reversed and magnets left 
     in standby mode.
    
    3. No detector problems.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Nov 01 2001 - 15:56:00 EST